OpenStudy (anonymous):

HELP ME Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). A line passes through point (-2, 5) and has a slope of 2/3. Points A(x, 3) and B(-2, y) lie on the line. The value of x is ______ and the value of y is______

OpenStudy (kamibug):

Here are some steps you could follow 1. Write the equation of the line in point-slope form\[y - y_1 = m(x - x_1)\] 2. Convert the equation to slope-intercept form\[y = mx + b\] 3. Graph the line (I attached the graph below.) 4. Find the points by looking for at what x value is on the line when y = 3 and what y value is on the line when x = -2.

OpenStudy (anonymous):

thanks I really appreciate it